OpenStudy (anonymous):

Find the slope of a line parallel to 5x + 2y = 6

OpenStudy (anonymous):

Will fan and medal ! PLEASE HELP !

OpenStudy (mrnood):

ok this is very simple IF you rewrite the equation in the form y=mx+b so FIRST SUbtract 5x from each side of the equation THEN Divide each side of the new equation by 2 Write your answer here and I will help with the final step...

OpenStudy (anonymous):

Okay so i got y=3-2.5x ?

OpenStudy (anonymous):

Or is it y-5x=-3? o.o

OpenStudy (mrnood):

ok so in this form you can know a lot about th eline the value before x is the slope the constant term is the intercept on the y axis SO now you know th eslope of the given line. The final step is tyo know that all lines PARALLEL to that line have the SAME slope (they just have a different intercept)

OpenStudy (anonymous):

Wait which one is the correct form?

OpenStudy (mrnood):

it has to be of the form y=mx~b so your first one is correct. The second is both incorrect AND of the wrong form!

OpenStudy (anonymous):

The first one i didn't even put into the correct form? I just left in the form it was in before.

OpenStudy (mrnood):

no - you have written it y=-2.5x +3 that is correct form (and correct equation)

OpenStudy (anonymous):

So would it now be written for parallel -2.5x - 3 ?

OpenStudy (mrnood):

It only asks you for the slope - you cannot know the full equation ALL parallel lines have the same slope